PLEASE BOLD THE FINAL ANSWERS AND MAKE THEM VISIBLE. I GIVE THUMBS UP

Mill Creek Corporation (MCC) currently has 450,000 shares of stock outstanding that sell for $90 per share. Assuming no market imperfections or tax effects exist, what will the share price be after:

a. MCC has a four-for-three stock split? (Do not round intermediate calculations. Round the final answer to 2 decimal places. Omit $ sign in your response.)

New share price           $  per share

b. MCC has a 10% stock dividend? (Do not round intermediate calculations. Round the final answer to 2 decimal places. Omit $ sign in your response.)

New share price           $  per share

c. MCC has a 47.5% stock dividend? (Do not round intermediate calculations. Round the final answer to 2 decimal places. Omit $ sign in your response.)

New share price           $  per share

d. MCC has a four-for-seven reverse stock split? (Do not round intermediate calculations. Round the final answer to 2 decimal places. Omit $ sign in your response.)

New share price           $  per share

Determine the new number of shares outstanding in parts (a) through (d). (Do not round intermediate calculations. Round the final answers to the nearest whole number.)

a. New shares outstanding. ______
b. New shares outstanding____
c. New shares outstanding____
d. New shares outstanding______
Answer #1

1.
=90*3/4=67.50

2.
=90/1.1=81.82

3.
=90/1.475=61.017

4.
=90*7/4=157.50

5.
=450000*4/3=600000

6.
=450000*1.1=495000

7.
=450000*1.475=663750

8.
=450000*4/7=257143